OpenStudy (anonymous):

simplify the rational expression (showing work). (m+5)/(m^(3)+10m^(2)+25m)

OpenStudy (anonymous):

look at it... m^3+10m^2+25m=m(m^2+10m+25)=m*(m+5)^2 so, (m+5)/(m*(m+5)^2)=1/(m*(m+5))